# Optimize Weight Rounding via Signed Gradient Descent for the Quantization of LLMs

### v4

@@ 4.1 Experimental Settings
V2-7B FP16 42.69 73.90 57.15 68.90 78.07 25.21 31.40 77.74 62.82 76.35 43.52 57.98
V2-13B FP16 52.86 76.77 60.04 72.14 79.05 25.95 35.20 80.55 65.34 79.38 48.38 61.42
V2-70B FP16 66.23 79.64 64.77 77.98 82.15 30.60 37.20 83.70 67.87 82.70 54.44 66.12
SignRound Hyperparameters.
Unless explicitly stated, the tuning process involved adjusting each block for 200 steps with a learning rate of $1/200$, i.e., $35e-$ 3, a batch size of 8, and linear learning rate decay. Additionally, we employed automatic mixed precision (AMP) to accelerate the tuning.
4.2 Comparing With Recent Methods
To ensure fair comparison as much as possible, we enabled act-order and true-sequential modes in GPTQ and also activated static_group in scenarios with group_size. The notation GPTQ+ indicates that we adjusted the random seed or data pre-processing to address issues related to the non-positive definite Hessian matrix or other issues. For OmniQuant(Shao et al., 2023), we adhere to the official settings, which include running for 20 epochs including W2G128 for saving time and disabling ’let’. We conducted calibration tests using sample sizes of 512 and 128, as well as a sample size of 512 with a batch size of 4. Our findings show that using a sample size of 512 typically results in comparable or slight higher performance for models less than or equal to 13B. Therefore, we present the results based on the sample size of 512. For 70B models, due the the NAN loss issue and to reduce the tuning cost of OmniQuant, we adopted 128 samples for calibration.
We present the summary results of Mistral-7B and LLAMAV2 in Table 1 , detailed results of W4G-1 in Table 2, and additional detailed results are provided in Appendix C due to space constraints. In summary, our approach demonstrated superior performance compared to GPTQ (Frantar et al., 2022), achieving scores of 30/32, AWQ (Lin et al., 2023) with 27/32, HQQ (Badri and Shaji, 2023) with 15/16, and OmniQuant (Shao et al., 2023) with a score of 29/32 across llamv1/llamav2/mistral-7b on various quantization settings, including W4G-1, W4G128, W3G128, and W2G128. These evaluations were based on the average accuracies of 11 zero-shot tasks.
It’s worth noting that as the bit depth decreases, the advantages of SignRound become more notable. For example, at W2G128, SignRound could yield absolute average accuracy improvements ranging from 6.91% to 33.22%.
Moreover, we can enhance the performance by optimizing the model’s hyperparameters from a selection of eight choices, denoted as ours*. These choices include steps (200, 1000), weight clip learning rate (1.0/steps, 2.0/steps), and the option to either enable or disable quantized inputs, which refers to utilizing the output from the previous quantized block or the previous original block.
4.3 Comparing with Rounding Methods
In this section, we conduct a comparative analysis between SignRound, FlexRound(Lee et al., 2023), and AdaRound(Nagel et al., 2020). Notably, there is no formal official implementation available for FlexRound and AdaRound for LLMs. Hence, we reference the code  [^3] https://openreview.net/forum?id=-tYCaP0phY_ for further details. However, it’s important to highlight that due to the lack of AMP support and other optimizations, the implementation is notably slow, especially when adhering to the official settings, which involve tuning 5000 steps, as presented in Table 9. Therefore, our comparison is limited to models of size 13B or smaller. We set the learning rate to 2e-4 for LLaMA-2-7b and Mistral-7B, and 1e-4 for LLaMA-13-7b to align with the official settings as closely as possible.
Table 3: Comparing with other rounding methods, the average accuracies ($\uparrow$) across 11 tasks (detailed in 4.1) for Mistral and LLaMA models at W4G-1 and W2G128.
@@ 4.3 Comparing with Rounding Methods
4.4 Ablation Studies
SignSGD versus Adam.
To validate the effectiveness of SignSGD, Table 4 compares it with the Adam optimizer Kingma and Ba (2014). SignSGD employs a fixed learning rate of $35e-$ 3 throughout all experiments, comprising 200 steps, with linear weight decay. For Adam, we explored learning rates ranging from $32.5e-$ 3 to $22e-$ 2. We choose to quantize models of 13B or less with W4G-1 due to the experiment’s cost. SignSGD demonstrated a distinct advantage in average accuracy metrics across 11 tasks, which demonstrate the unique advantage of signed gradient descent in this scenario.
Table 4: Adam optimizer with various learning rates vs. SignSGD optimizer. The average accuracies($\uparrow$) across 11 tasks (detailed in 4.1) for Mistral and LLaMA models at W4G-1 and W2G128.
@@ 4.4 Ablation Studies
Round and Weigh Clip Tuning.
Table 5: Ablation study of Round tuning and weight clip tuning. The average accuracies($\uparrow$) across 11 tasks(detailed in 4.1) for Mistral and LLaMA models at W4G-1 and W2G128.
@@
Hypeparameters Sensitivity.
Table 6: Ablation study of of hyperparameter sensitivity. The average accuracies($\uparrow$) across 11 tasks(detailed in 4.1) for LLaMA models at W4G-1.
@@ 4.5 Generalization to Other Models
To assess the generalization of our method on LLMs, we evaluate SignRound on various mainstream LLMs such as Gemma (Team et al., 2024), Phi (Li et al., 2023b), Mistral (Jiang et al., 2023), Mixtral (Jiang et al., 2024) and Llama3 (Touvron et al., 2024). Table 7 demonstrated that all int4 models met the requirement of maintaining accuracy drop within 1% of the FP16 or BF16 accuracy by employing 1000 tuning steps and model wise hype-parameters among 4 choices detailed in section 4.1. Additionally, one-third of the models exhibited a slight improvement in accuracy compared to the baseline. Please note that the generalization experiments utilize an updated version (0.4.0+) of lm-eval-harness (Gao et al., 2023) and real quantized model, which may result in minor discrepancies compared to other benchmark data.
Table 7: Accuracies($\uparrow$) across 11 tasks(0-shot) with 1000 steps for LLMs at W4G128
Model Method Mmlu Lamb. Hella. Wino. Piqa Truth. Open. Boolq RTE ARC-e ARC-c. Avg. Vari. %
gemma-7b FP16 61.62 67.51 60.47 73.24 79.43 30.97 33.20 82.78 65.34 81.78 49.91 62.39 -
Ours 61.47 72.04 59.03 75.14 79.49 30.11 34.00 82.69 70.76 79.59 49.40 63.07 1.09%
gemma-7b-it FP16 50.29 60.35 56.20 67.96 77.09 30.48 37.40 81.38 78.70 75.25 47.27 60.22 -
Ours 49.93 62.86 55.64 67.88 77.31 30.35 37.00 81.44 78.70 75.08 45.73 60.17 -0.08%
Llama-2-7b-chat FP16 46.40 71.05 57.80 66.38 76.39 30.23 33.40 79.76 69.68 73.82 44.20 59.01 -
Ours 45.45 70.37 57.06 66.14 76.33 30.35 32.60 80.64 72.92 73.36 43.52 58.97 -0.07%
Llama-3-8B-Instruct BF16 63.86 71.82 57.69 71.43 78.67 36.23 34.00 82.97 67.51 81.52 52.99 63.52 -
Ours 63.06 72.00 56.99 72.38 77.97 35.37 33.00 83.09 68.59 80.89 51.02 63.12 -0.63%
Mistral-7B BF16 59.61 75.61 61.28 74.43 80.79 28.03 32.80 83.73 66.43 80.85 50.09 63.06 -
Ours 58.80 75.51 60.79 74.51 80.14 28.89 33.00 82.78 69.68 80.60 49.15 63.08 0.03%
Mixtral-8x7B BF16 68.02 78.27 64.90 76.48 82.48 34.27 35.40 85.23 70.76 84.30 56.66 66.98 -
Ours 66.93 78.25 64.59 75.14 82.10 32.19 35.60 84.74 69.31 84.30 56.48 66.33 -0.97%
Mixtral-8x7B-Instruct BF16 68.85 77.18 67.67 76.87 83.51 49.69 36.80 88.50 71.84 86.99 62.20 70.00 -
Ours 68.24 77.90 67.45 77.19 83.35 48.84 37.20 87.83 70.04 87.12 62.29 69.77 -0.33%
Phi-3-mini-4k-instruct FP16 68.20 68.14 60.64 73.88 80.14 39.53 38.40 86.24 78.34 83.42 55.63 66.93
Ours 67.28 66.78 59.69 73.17 79.43 37.58 39.00 85.81 77.98 83.33 56.48 66.33 -0.90%
5 Conclusions and Limitations
In this paper, we introduce SignRound, an efficient and concise approach for optimizing weight rounding in the quantization of large language models. SignRound employs signed gradient descent for tuning rounding value and weight clipping in 200 steps, completing the quantization of LLAMA-V2-70B in approximately 2.5 hours. Our extensive experiments show that SignRound outperforms other quantization methods across various models and weight bits, maintaining fixed hyperparameter settings. Additionally, SignRound shows promising generation capabilities in recent models and achieves enhanced performance through model-specific hyperparameter tuning.
Despite these significant advantages, we still observed a noticeable gap in the accuracy performance of ultra-low bit quantization compared to the original model, such as 2-bit quantization. This challenge could potentially be addressed by exploring non-uniform quantization and mixed-precision quantization, which we leave for future work.
@@ Appendix A Quantization Cost
Table 8 compares the quantization costs of different methods, with all measurements conducted on a single NVIDIA A100 GPU with 80GB of memory. We ensure each evaluation process exclusively occupies one GPU, but CPU and other resources may be shared among different processes due to limited resources. For SignRound, we disabled low_gpu_mem_usage in our implementation to achieve faster tuning, albeit with higher memory usage. Despite this, LLaMAV2-70B was still able to run on an A100 GPU with 80GB of memory. Although HQQ is exceptionally fast, our methods outperform others in terms of speed. Table 9 also compares the costs between FlexRound, Adaptive Round, and our method.
Table 8: Quantization cost in seconds at W4G-1 for LLaMA2. Align with the accuracy experiments, OmniQuant 70b is tested with 128 calibration samples, while all the others are tested with 512 samples.
Model GPTQ AWQ HQQ OmniQuant Ours
LLaMAV2-7B 1821 1328 19 10255 1041
LLaMAV2-13B 3266 2630 30 18186 1918
LLaMAV2-70B 18517 13586 119 35694 9116
@@
Method FlexRound AdaRound Ours
Mistral-7B-V0.1 9369 9332 1045
LLaMAV2-7B 9628 9701 1041
LLaMAV2-13B 17583 17865 1918
Appendix B View of distribution of tuned parameters
Figure 2 illustrates the distribution of the magnitudes of V in Eq.3 and $\alpha$, $\beta$ in Eq. 4 for Mistral-7B-v0.1 and LLaMA-2-7B at W4G-1. The results indicate that the distribution is flat for most layers, except for a few layers at the beginning and the end.
@@ Appendix B View of distribution of tuned parameters
Figure 2: The distribution of the magnitude of V in Eq. 3 and $\alpha$, $\beta$ in Eq. 4 for Mistral-7B-v0.1 and LLaMA-2-7B at W4G-1, each color in the distribution represents a specific layer index in the models, with blue indicating shallow layers closer to the data layer, and red representing deeper layers.
V1-7B FP16 32.74 73.53 56.94 70.01 78.67 22.03 34.60 75.08 66.43 75.25 41.81 57.01
V1-13B FP16 44.21 76.21 59.92 72.77 79.16 25.70 33.20 77.89 70.76 77.40 46.42 60.33
V1-30B FP16 55.14 77.55 63.33 75.85 81.12 28.27 36.00 82.78 66.79 80.39 52.90 63.65
V1-65B FP16 59.79 79.12 64.53 77.35 81.23 27.91 38.00 84.86 69.68 81.36 52.82 65.15
FP16 61.35 75.68 61.27 74.03 80.79 28.03 32.80 83.67 67.51 80.81 50.34 63.30
FP16 42.69 73.90 57.15 68.90 78.07 25.21 31.40 77.74 62.82 76.35 43.52 57.98
FP16 52.86 76.77 60.04 72.14 79.05 25.95 35.20 80.55 65.34 79.38 48.38 61.42
FP16 66.23 79.64 64.77 77.98 82.15 30.60 37.20 83.70 67.87 82.70 54.44 66.12
V1-7B FP16 32.74 73.53 56.94 70.01 78.67 22.03 34.60 75.08 66.43 75.25 41.81 57.01
V1-13B FP16 44.21 76.21 59.92 72.77 79.16 25.70 33.20 77.89 70.76 77.40 46.42 60.33
V1-30B FP16 55.14 77.55 63.33 75.85 81.12 28.27 36.00 82.78 66.79 80.39 52.90 63.65
V1-65B FP16 59.79 79.12 64.53 77.35 81.23 27.91 38.00 84.86 69.68 81.36 52.82 65.15
FP16 61.35 75.68 61.27 74.03 80.79 28.03 32.80 83.67 67.51 80.81 50.34 63.30
FP16 42.69 73.90 57.15 68.90 78.07 25.21 31.40 77.74 62.82 76.35 43.52 57.98
FP16 52.86 76.77 60.04 72.14 79.05 25.95 35.20 80.55 65.34 79.38 48.38 61.42
FP16 66.23 79.64 64.77 77.98 82.15 30.60 37.20 83.70 67.87 82.70 54.44 66.12
V1-7B FP16 32.74 73.53 56.94 70.01 78.67 22.03 34.60 75.08 66.43 75.25 41.81 57.01
V1-13B FP16 44.21 76.21 59.92 72.77 79.16 25.70 33.20 77.89 70.76 77.40 46.42 60.33
V1-30B FP16 55.14 77.55 63.33 75.85 81.12 28.27 36.00 82.78 66.79 80.39 52.90 63.65
V1-65B FP16 59.79 79.12 64.53 77.35 81.23 27.91 38.00 84.86 69.68 81.36 52.82 65.15
FP16 61.35 75.68 61.27 74.03 80.79 28.03 32.80 83.67 67.51 80.81 50.34 63.30
FP16 42.69 73.90 57.15 68.90 78.07 25.21 31.40 77.74 62.82 76.35 43.52 57.98
FP16 52.86 76.77 60.04 72.14 79.05 25.95 35.20 80.55 65.34 79.38 48.38 61.42
FP16 66.23 79.64 64.77 77.98 82.15 30.60 37.20 83.70 67.87 82.70 54.44 66.12
V1-7B FP16 32.74 73.53 56.94 70.01 78.67 22.03 34.60 75.08 66.43 75.25 41.81 57.01
V1-13B FP16 44.21 76.21 59.92 72.77 79.16 25.70 33.20 77.89 70.76 77.40 46.42 60.33
V1-30B FP16 55.14 77.55 63.33 75.85 81.12 28.27 36.00 82.78 66.79 80.39 52.90 63.65
V1-65B FP16 59.79 79.12 64.53 77.35 81.23 27.91 38.00 84.86 69.68 81.36 52.82 65.15
Table 14: Perplexity(PPL) ($\downarrow$) of Wikitext2, PTB , C4 and Wikitext tasks for LLaMA and Mistral models. we follow the source code of GPTQ for wikitext2, PTB and C4 PPL evaluation, while for wikitext, we adopt lm-eval-harness [Gao et al., 2023]. NAN indicates not a number, while OOM denotes out of memory.

### v5

@@ 4.1 Experimental Settings
16 bits 61.35 75.68 61.27 74.03 80.79 28.03 32.80 83.67 67.51 80.81 50.34 63.30
W2G128 AWQ 25.38 0.00 25.71 52.01 51.58 23.99 17.60 37.83 47.29 26.98 22.27 30.06
Ours* 43.72 59.75 51.87 64.25 75.14 24.72 23.60 75.78 55.23 71.80 37.20 53.01
Table 2: Detailed accuracies($\uparrow$) across 11 tasks(0-shot) of Mistral models at W4G-1 and W2G128. ’Our*’ denotes the highest accuracy achieved among the 8 hyperparameter choices, outlined in Section 4.2. Appendix C provides more detailed data.
@@
Table 3: Comparing with some other rounding methods, the average accuracies ($\uparrow$) across 11 tasks (detailed in Section 4.1) for Mistral and LLaMA models at W4G-1 and W2G128.
Evaluation and Tasks.
We evaluate multiple language tasks to address the task-agnostic setting. Specifically, we present the average accuracy results for 11 zero-shot tasks, including HellaSwag (Zellers et al., 2019), WinoGrande (Sakaguchi et al., 2021), PIQA (Bisk et al., 2020), LAMBADA (Paperno et al., 2016), TruthfulQA (Lin et al., 2022), OpenBookQA (Mihaylov et al., 2018), BoolQ (Clark et al., 2019), RTE (Dagan et al., 2010), ARC-Easy, ARC-Challenge (Clark et al., 2018), and MMLU (Hendrycks et al., 2020). We use lm-eval-harness (Gao et al., 2023) for all the above tasks. Furthermore, we complement our evaluation with perplexity (PPL) analysis on Wikitext2 (Merity et al., 2016), PTB (Marcus et al., 1993), and C4 (Raffel et al., 2020), following the implementation [^1] https://github.com/IST-DASLab/gptq of GPTQ and Wikitext2 (Merity et al., 2016) using lm-eval-harness (Gao et al., 2023). However, we argue that perplexity is notably influenced by outliers, as illustrated in Table 14 for different algorithms. This susceptibility likely arises from the mathematical expression $\mathrm{PPL}(X)=\exp\left(-\frac{1}{t}\sum_{i=1}^{t}\log p_{\theta}(x_{i}|x_{<i})\right)$, where assigning a low probability to even one token can significantly inflate the perplexity score. Consequently, we prioritize the accuracy of the 11 tasks mentioned above as the primary metric, with perplexity data serving as supplementary reference.
Quantization Configurations.
In alignment with GPTQ (Frantar et al., 2022), our focus is specifically on weight-only quantization, targeting the linear layers within transformer blocks. Layers such as the embedding layer and typically the last linear layer like ’lm-head’ are excluded from the quantization process. Our evaluation primarily centers on W4G-1, W4G128, W3G128 and W2G128 configurations, where W4 indicates quantizing weights with 4 bits and G represents finer-grained grouping as described in (Park et al., 2022; Frantar et al., 2022). We adopt asymmetric quantization. To mitigate overfitting on the WikiText and C4 datasets, for all the methods that need calibration, we randomly select 512 calibration samples with the same seed from the readily available pile-10k dataset  [^2] https://huggingface.co/datasets/NeelNanda/pile-10k , which comprises the first 10k samples from pile (Gao et al., 2020). We used a sequence length of 2048 for calibration, while for other methods, we adhere to their official settings.
Large Language Models.
SignRound Hyperparameters.
Unless explicitly stated, the tuning process involved adjusting each block for 200 steps with a learning rate of $5\times 10^{-3}$, a batch size of 8, and linear learning rate decay. Additionally, we employed automatic mixed precision (AMP) to accelerate the tuning.
4.2 Comparing With Recent Methods
To ensure fair comparison as much as possible, we enabled act-order and true-sequential in GPTQ and also activated static_group in scenarios with group_size. The notation GPTQ+ indicates that we adjusted the random seed or data pre-processing to address issues related to the non-positive definite Hessian matrix or other issues. For OmniQuant(Shao et al., 2023), we adhere to the official settings, which include running for 20 epochs including W2G128 for saving time and disabling ’let’. We conducted calibration tests using sample sizes of 512 and 128, as well as a sample size of 512 with a batch size of 4. Our findings show that using a sample size of 512 typically results in comparable or slightly higher performance for models less than or equal to 13B. Therefore, we present the results based on the sample size of 512. For 70B models, due the the Not a Number (NAN) loss issue and to reduce the tuning cost of OmniQuant, we adopted 128 samples for calibration.
@@ 4.2 Comparing With Recent Methods
Table 4: Comparison of Adam optimizer with various learning rates against the SignSGD optimizer.. The average accuracies($\uparrow$) across 11 tasks (detailed in Section 4.1) for Mistral and LLaMA models at W4G-1 and W2G128.
@@
Table 5: Ablation study of round tuning and weight clip tuning. The average accuracies($\uparrow$) across 11 tasks(detailed in Section 4.1) for Mistral and LLaMA models at W4G-1 and W2G128.
We present the summary results of Mistral-7B and LLaMAV2 in Table 1, detailed results of Mistral-7B in Table 2, and additional detailed results are provided in Appendix C due to space constraints. In summary, our approach demonstrated superior performance compared to GPTQ (Frantar et al., 2022), achieving scores of 30/32, AWQ (Lin et al., 2023) with 27/32, HQQ (Badri and Shaji, 2023) with 15/16, and OmniQuant (Shao et al., 2023) with a score of 29/32 across LLaMAV1/LLaMAV2/Mistral-7B on various quantization settings, including W4G-1, W4G128, W3G128, and W2G128. These evaluations were based on the average accuracies of 11 zero-shot tasks.
It’s worth noting that as the bit depth decreases, the advantages of SignRound become more notable. For example, as shown in Table 2, SignRound could yield absolute average accuracy improvements ranging from 6.91% to 33.22% at W2G128.
Moreover, we can enhance the performance by tuning the model’s hyperparameters from a selection of eight choices, denoted as ours*. These choices include steps (200, 1000), weight clip learning rate (1.0/steps, 2.0/steps), and the option to either enable or disable quantized inputs, which refers to utilizing the output from the previous quantized block or the previous original block.
4.3 Comparing with Rounding Methods
In this section, we conduct a comparative analysis between SignRound, FlexRound(Lee et al., 2023), and AdaRound(Nagel et al., 2020). Notably, during the experiment, there is no formal official implementation available for FlexRound and AdaRound for LLMs. Hence, we reference the implementations  [^3] https://openreview.net/forum?id=-tYCaP0phY_  [^4] https://github.com/quic/aimet for further details. However, it’s important to highlight that due to the lack of AMP support and other optimizations, the implementation is notably slow, especially when adhering to the official settings, which involve tuning 5000 steps, as presented in Table 9. Therefore, our comparison is limited to models of size 13B or smaller. We set the learning rate to 2e-4 for LLaMA-v2-7b and Mistral-7B, and 1e-4 for LLaMA-v2-13b to align with the official settings as closely as possible. As shown in Table 3, SignRound achieves better results in just 200 steps compared to the 5000 steps required by other rounding methods.
4.4 Ablation Studies
SignSGD versus Adam.
To validate the effectiveness of SignSGD, Table 4 compares it with the Adam optimizer Kingma and Ba (2014). SignSGD employs a fixed learning rate of 5e-3 throughout all experiments, comprising 200 steps, with linear weight decay. For Adam, we explored learning rates ranging from 2.5e-3 to 2e-2. We choose to quantize models of 13B or less with W4G-1 due to the experiment’s cost. SignSGD demonstrated a distinct advantage in average accuracy metrics across 11 tasks, which demonstrate the unique advantage of signed gradient descent in this scenario.
Round and Weigh Clip Tuning.
@@ 4.4 Ablation Studies
Table 6: Ablation study of hyperparameter sensitivity. The average accuracies($\uparrow$) across 11 tasks(detailed in Section 4.1) for LLaMA models at W4G-1.
Hyperparameters Sensitivity.
Model Method Average Acc Variation %
Gemma-2b BF16 53.69 -
Ours 53.40 -0.54%
Llama-2-7b-chat-hf FP16 59.01 -
Ours 58.97 -0.07%
Llama-3-8B-Instruct BF16 63.52 -
Ours 63.12 -0.63%
Mistral-7B-Instruct-v0.2 BF16 66.47 -
Ours 66.21 -0.39%
Mixtral-8x7B BF16 66.98 -
Ours 66.33 -0.97%
Mixtral-8x7B-Instruct BF16 70.00 -
Ours 69.77 -0.33%
Phi-3-mini-4k-instruct BF16 66.62 -
Ours 66.33 -0.44%
Table 7: The average accuracies($\uparrow$) across 11 tasks(detailed in Section 4.1)) with 1000 steps for LLMs at W4G128. Table 15 provides the detailed data.
@@ 4.5 Generalization to Other Models
To assess the generalization of our method on LLMs, we evaluate SignRound on various mainstream LLMs such as Gemma (Team et al., 2024), Phi (Li et al., 2023b), Mistral (Jiang et al., 2023), Mixtral (Jiang et al., 2024) and Llama3 (Touvron et al., 2024). Table 7 demonstrated that all int4 models maintained an accuracy drop within 1% of FP16 or BF16 accuracy by employing 1000 tuning steps and model wise hyperparameters among 4 choices detailed in Section 4.1. The detailed results are provided in Appendix C. Notably, the generalization experiments utilized an updated version (0.4.0+) of lm-eval-harness (Gao et al., 2023) and real quantized models, which may result in minor discrepancies compared to other benchmark data.
5 Conclusions
In this paper, we introduce SignRound, an efficient and concise approach for optimizing weight rounding in the quantization of large language models. SignRound employs signed gradient descent for tuning rounding value and weight clipping in 200 steps, completing the quantization of LLAMA-V2-70B in approximately 2.5 hours. Our extensive experiments show that SignRound outperforms other quantization methods across various models and weight bits in the majority of scenarios. Additionally, SignRound shows promising generation capabilities in recent models and achieves enhanced performance through model-specific hyperparameter tuning.
6 Limitations
Despite the advantages, we observed a noticeable gap in accuracy performance for ultra-low bit quantization, particularly with 2-bit quantization, compared to the original model. This challenge could potentially be addressed by exploring non-uniform quantization and mixed-precision quantization, which we leave for future work.
7 Ethics Statement
Our research aims to advance knowledge in LLM quantization. SignRound utilizes open-source models and publicly available datasets, and is not tied to particular applications, requiring only minimal fine-tuning steps on the original models. This ensures that the technical details of our method carry no potential ethical implications. We acknowledge the contributions of the creators and maintainers of these resources and provide citations to the original sources.
@@ Appendix A Quantization Cost
Table 8 compares the quantization costs of different methods, with all measurements conducted on a single NVIDIA A100 GPU with 80GB of memory. We ensure each evaluation process exclusively occupies one GPU, but CPU and other resources may be shared among different processes due to limited resources. For SignRound, we disabled low_gpu_mem_usage in our implementation to achieve faster tuning, albeit with higher memory usage. Despite this, LlaMA-2-70B was still able to run on an A100 GPU with 80GB of memory. Although HQQ is exceptionally fast, our methods outperform others in terms of speed. Table 9 also compares the costs between FlexRound, Adaptive Round, and our method.
Model GPTQ AWQ HQQ Omni. Ours
Llama-2-7B 1821 1328 19 10255 1041
Llama-2-13B 3266 2630 30 18186 1918
Llama-2-70B 18517 13586 119 35694 9116
Table 8: Quantization cost in seconds at W4G-1 for LLaMAV2. Align with the accuracy experiments, OmniQuant 70b is tested with 128 calibration samples, while all the others are tested with 512 samples.
Appendix B View of distribution of tuned parameters
Figure 2 illustrates the distribution of the magnitudes of $\mathbf{V}$ in Eq.3 and $\alpha$, $\beta$ in Eq. 4 for Mistral-7B-v0.1 and Llama-2-7B at W4G-1. The results indicate that the distribution is flat for most layers, except for a few layers at the beginning and the end.
Model FlexRound AdaRound Ours
Mistral-7B-V0.1 9369 9332 1045
Llama-2-7B 9628 9701 1041
Llama-2-13B 17583 17865 1918
@@ Appendix B View of distribution of tuned parameters
Mistral-7B 16 bits 61.35 75.68 61.27 74.03 80.79 28.03 32.80 83.67 67.51 80.81 50.34 63.30
V2-7B 16 bits 42.69 73.90 57.15 68.90 78.07 25.21 31.40 77.74 62.82 76.35 43.52 57.98
V2-13B 16 bits 52.86 76.77 60.04 72.14 79.05 25.95 35.20 80.55 65.34 79.38 48.38 61.42
V2-70B 16 bits 66.23 79.64 64.77 77.98 82.15 30.60 37.20 83.70 67.87 82.70 54.44 66.12
V1-7B 16 bits 32.74 73.53 56.94 70.01 78.67 22.03 34.60 75.08 66.43 75.25 41.81 57.01
V1-13B 16 bits 44.21 76.21 59.92 72.77 79.16 25.70 33.20 77.89 70.76 77.40 46.42 60.33
V1-30B 16 bits 55.14 77.55 63.33 75.85 81.12 28.27 36.00 82.78 66.79 80.39 52.90 63.65
V1-65B 16 bits 59.79 79.12 64.53 77.35 81.23 27.91 38.00 84.86 69.68 81.36 52.82 65.15
16 bits 61.35 75.68 61.27 74.03 80.79 28.03 32.80 83.67 67.51 80.81 50.34 63.30
16 bits 42.69 73.90 57.15 68.90 78.07 25.21 31.40 77.74 62.82 76.35 43.52 57.98
16 bits 52.86 76.77 60.04 72.14 79.05 25.95 35.20 80.55 65.34 79.38 48.38 61.42
16 bits 66.23 79.64 64.77 77.98 82.15 30.60 37.20 83.70 67.87 82.70 54.44 66.12
V1-7B 16 bits 32.74 73.53 56.94 70.01 78.67 22.03 34.60 75.08 66.43 75.25 41.81 57.01
V1-13B 16 bits 44.21 76.21 59.92 72.77 79.16 25.70 33.20 77.89 70.76 77.40 46.42 60.33
V1-30B 16 bits 55.14 77.55 63.33 75.85 81.12 28.27 36.00 82.78 66.79 80.39 52.90 63.65
V1-65B 16 bits 59.79 79.12 64.53 77.35 81.23 27.91 38.00 84.86 69.68 81.36 52.82 65.15
16 bits 61.35 75.68 61.27 74.03 80.79 28.03 32.80 83.67 67.51 80.81 50.34 63.30
16 bits 42.69 73.90 57.15 68.90 78.07 25.21 31.40 77.74 62.82 76.35 43.52 57.98
16 bits 52.86 76.77 60.04 72.14 79.05 25.95 35.20 80.55 65.34 79.38 48.38 61.42
16 bits 66.23 79.64 64.77 77.98 82.15 30.60 37.20 83.70 67.87 82.70 54.44 66.12
V1-7B 16 bits 32.74 73.53 56.94 70.01 78.67 22.03 34.60 75.08 66.43 75.25 41.81 57.01
V1-13B 16 bits 44.21 76.21 59.92 72.77 79.16 25.70 33.20 77.89 70.76 77.40 46.42 60.33
V1-30B 16 bits 55.14 77.55 63.33 75.85 81.12 28.27 36.00 82.78 66.79 80.39 52.90 63.65
V1-65B 16 bits 59.79 79.12 64.53 77.35 81.23 27.91 38.00 84.86 69.68 81.36 52.82 65.15
16 bits 61.35 75.68 61.27 74.03 80.79 28.03 32.80 83.67 67.51 80.81 50.34 63.30
16 bits 42.69 73.90 57.15 68.90 78.07 25.21 31.40 77.74 62.82 76.35 43.52 57.98
16 bits 52.86 76.77 60.04 72.14 79.05 25.95 35.20 80.55 65.34 79.38 48.38 61.42
16 bits 66.23 79.64 64.77 77.98 82.15 30.60 37.20 83.70 67.87 82.70 54.44 66.12
V1-7B 16 bits 32.74 73.53 56.94 70.01 78.67 22.03 34.60 75.08 66.43 75.25 41.81 57.01
V1-13B 16 bits 44.21 76.21 59.92 72.77 79.16 25.70 33.20 77.89 70.76 77.40 46.42 60.33
V1-30B 16 bits 55.14 77.55 63.33 75.85 81.12 28.27 36.00 82.78 66.79 80.39 52.90 63.65
V1-65B 16 bits 59.79 79.12 64.53 77.35 81.23 27.91 38.00 84.86 69.68 81.36 52.82 65.15
@@
Figure 2: The distribution of the magnitude of $\mathbf{V}$ in Eq. 3 and $\alpha$, $\beta$ in Eq. 4 for Mistral-7B-v0.1 and Llama-2-7B at W4G-1, each color in the distribution represents a specific layer index in the models, with blue indicating shallow layers closer to the data layer, and red representing deeper layers.